OpenStudy (anonymous):

Simplify:

OpenStudy (anonymous):

\[(5+2\sqrt{7})^{2}\]

OpenStudy (saifoo.khan):

\[(5+2\sqrt{7})(5+2\sqrt{7})\]Use FOIL to solve.

OpenStudy (anonymous):

How do I use FOIL again? Its been so long.

OpenStudy (jamesj):

\[(a+b)(c+d) = ac + ad + bc + bd \]

OpenStudy (anonymous):

That is so confusing.. I don't understand it at all. Things have to be shown to me step by step..

OpenStudy (jamesj):

\[ (a+b)(c+d)\] \[ = a(c+d) + b(c+d) \] \[ = ac + ad + bc + bd \]

OpenStudy (anonymous):

That just confused me even more.. thanks.

OpenStudy (jamesj):

what's a(c + d) equal to?

OpenStudy (anonymous):

I don't have any clue.. is there any easier way you could show me thiss??

OpenStudy (jamesj):

Err ... what's 4*(10 + 11) equal to?

OpenStudy (jamesj):

or just 4(10 + 11)

OpenStudy (anonymous):

84..

OpenStudy (jamesj):

is it true that 4(10 + 11) = 4*10 + 4*11 ?

OpenStudy (anonymous):

Yes.

OpenStudy (jamesj):

So is true that 4(c + d) = 4c + 4d ?

OpenStudy (anonymous):

Yes.

OpenStudy (jamesj):

And is true that a(c + d) = ac + ad ?

OpenStudy (anonymous):

Yes..

OpenStudy (jamesj):

and b(c + d) = bc + bd?

OpenStudy (anonymous):

Mhmmm.

OpenStudy (saifoo.khan):